OpenStudy (anonymous):

Find a positive number that equals 90 when added to its square

OpenStudy (anonymous):

Check for 9..

OpenStudy (anonymous):

x+x^2=90

OpenStudy (anonymous):

is that the answer? @myko

OpenStudy (anonymous):

no

OpenStudy (anonymous):

solve this equation

OpenStudy (anonymous):

x^2+x-90=0

OpenStudy (anonymous):

Factors can be 10 and 9..

OpenStudy (anonymous):

like @waterineyes said, put it in this form (x+10)(x-9)=0 so x=-10, and x=9 since the problem sais it has to be positive, it means x=9 is your answer